# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v3

@@ Appendix C Processing times
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
@@
AeT 0.12 0.40 3.67 0.14 0.13 6.86 10.06
TAeT 0.19 0.40 3.62 0.15 0.14 5.99 10.15
JeT 0.07 0.18 3.22 0.07 0.07 4.05 9.81
TJeT 0.07 0.24 2.74 0.08 0.07 3.90 8.94
AeJeT 0.11 0.44 3.63 0.15 0.14 4.32 9.97
JeAeT 0.08 0.23 2.99 0.09 0.08 3.89 10.31
TAeTJeT 0.12 0.42 3.71 0.16 0.15 5.04 10.46
TJeTAeT 0.09 0.27 2.86 0.10 0.10 4.11 9.43
AJeT 0.12 0.45 3.61 0.15 0.15 4.20 10.17
Greedy 0.13 0.45 3.93 0.16 0.16 4.91 10.18
Appendix D Performance by Language
@@
D.1 Chinese
D.1.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.87 0.68 0.86 0.82 0.82 0.72 0.74
AeJeT 0.87 0.68 0.86 0.82 0.82 0.72 0.75
AeT 0.87 0.72 0.87 0.84 0.84 0.77 0.72
Greedy 0.87 0.75 0.86 0.84 0.84 0.76 0.76
JeAeT 0.88 0.65 0.86 0.82 0.82 0.71 0.76
JeT 0.90 0.66 0.87 0.83 0.83 0.72 0.76
@@ D.1.1 Precision
TAeT 0.87 0.75 0.86 0.84 0.84 0.77 0.74
TAeTJeT 0.87 0.75 0.87 0.84 0.84 0.76 0.77
TJeT 0.99 0.76 0.98 0.94 0.94 0.81 0.81
TJeTAeT 0.97 0.76 0.96 0.92 0.92 0.80 0.82
D.1.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.73 0.76 0.67 0.75 0.75 0.80 0.38
AeJeT 0.73 0.76 0.67 0.75 0.75 0.80 0.40
AeT 0.76 0.76 0.69 0.77 0.77 0.81 0.42
Greedy 0.76 0.71 0.70 0.77 0.77 0.82 0.37
JeAeT 0.64 0.73 0.61 0.67 0.67 0.73 0.32
JeT 0.63 0.73 0.60 0.66 0.66 0.73 0.33
@@ D.1.2 Recall
TAeT 0.75 0.70 0.70 0.77 0.77 0.81 0.39
TAeTJeT 0.76 0.70 0.70 0.77 0.77 0.82 0.38
TJeT 0.70 0.70 0.63 0.70 0.70 0.78 0.29
TJeTAeT 0.71 0.70 0.64 0.71 0.71 0.78 0.30
D.2 English
D.2.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.69 0.76 0.67 0.73 0.73 0.73 0.70
AeJeT 0.69 0.76 0.67 0.73 0.73 0.73 0.70
AeT 0.73 0.78 0.70 0.77 0.77 0.78 0.73
Greedy 0.72 0.78 0.70 0.77 0.77 0.79 0.72
JeAeT 0.64 0.68 0.64 0.69 0.69 0.67 0.65
JeT 0.64 0.69 0.64 0.69 0.69 0.67 0.65
@@ D.2.1 Precision
TAeT 0.73 0.78 0.70 0.77 0.77 0.78 0.74
TAeTJeT 0.72 0.78 0.70 0.77 0.77 0.78 0.73
TJeT 0.70 0.75 0.69 0.75 0.75 0.76 0.70
TJeTAeT 0.70 0.75 0.69 0.75 0.75 0.76 0.71
D.2.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.79 0.68 0.88 0.86 0.86 0.80 0.73
AeJeT 0.79 0.68 0.88 0.86 0.86 0.80 0.73
AeT 0.78 0.69 0.88 0.85 0.85 0.79 0.72
Greedy 0.81 0.71 0.89 0.86 0.86 0.81 0.76
JeAeT 0.74 0.55 0.87 0.85 0.85 0.76 0.66
JeT 0.75 0.55 0.88 0.86 0.86 0.77 0.66
@@ D.2.2 Recall
TAeT 0.78 0.69 0.88 0.86 0.86 0.80 0.73
TAeTJeT 0.81 0.71 0.89 0.86 0.86 0.81 0.76
TJeT 0.85 0.73 0.93 0.91 0.91 0.83 0.79
TJeTAeT 0.84 0.73 0.92 0.90 0.90 0.83 0.79
D.3 French
D.3.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.88 0.81 0.92 0.84 0.84 0.77 0.74
AeJeT 0.88 0.81 0.92 0.84 0.84 0.77 0.75
AeT 0.89 0.83 0.91 0.86 0.86 0.79 0.76
Greedy 0.90 0.84 0.92 0.86 0.86 0.82 0.76
JeAeT 0.93 0.82 0.94 0.87 0.87 0.72 0.69
JeT 0.94 0.83 0.94 0.88 0.88 0.72 0.70
@@ D.3.1 Precision
TAeT 0.88 0.84 0.92 0.86 0.86 0.81 0.76
TAeTJeT 0.89 0.84 0.91 0.86 0.86 0.82 0.77
TJeT 0.96 0.89 0.95 0.92 0.92 0.86 0.82
TJeTAeT 0.95 0.89 0.95 0.91 0.91 0.86 0.80
D.3.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.75 0.80 0.77 0.88 0.88 0.85 0.85
AeJeT 0.75 0.80 0.77 0.88 0.88 0.85 0.85
AeT 0.78 0.85 0.77 0.90 0.90 0.89 0.88
Greedy 0.77 0.85 0.78 0.91 0.91 0.89 0.89
JeAeT 0.70 0.76 0.77 0.86 0.86 0.84 0.84
JeT 0.69 0.75 0.77 0.86 0.86 0.84 0.84
@@ D.3.2 Recall
TAeT 0.78 0.85 0.77 0.90 0.90 0.89 0.89
TAeTJeT 0.77 0.85 0.78 0.91 0.91 0.89 0.89
TJeT 0.75 0.84 0.79 0.90 0.90 0.89 0.90
TJeTAeT 0.76 0.85 0.79 0.91 0.91 0.89 0.89
D.4 German
D.4.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.88 0.82 0.87 0.86 0.86 0.79 0.85
AeJeT 0.87 0.82 0.87 0.86 0.86 0.79 0.84
AeT 0.87 0.82 0.88 0.87 0.87 0.80 0.84
Greedy 0.88 0.83 0.89 0.87 0.87 0.81 0.86
JeAeT 0.87 0.81 0.86 0.84 0.84 0.78 0.83
JeT 0.88 0.82 0.86 0.84 0.84 0.78 0.83
@@ D.4.1 Precision
TAeT 0.88 0.83 0.89 0.87 0.87 0.80 0.85
TAeTJeT 0.88 0.84 0.89 0.87 0.87 0.81 0.86
TJeT 0.93 0.86 0.92 0.91 0.91 0.83 0.89
TJeTAeT 0.92 0.85 0.91 0.90 0.90 0.83 0.89
D.4.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.73 0.79 0.81 0.85 0.85 0.83 0.75
AeJeT 0.72 0.78 0.81 0.85 0.85 0.83 0.75
AeT 0.76 0.81 0.84 0.89 0.89 0.87 0.78
Greedy 0.76 0.84 0.85 0.89 0.89 0.88 0.79
JeAeT 0.70 0.76 0.80 0.84 0.84 0.82 0.74
JeT 0.67 0.73 0.77 0.82 0.82 0.80 0.70
@@ D.4.2 Recall
TAeT 0.76 0.82 0.84 0.89 0.89 0.87 0.79
TAeTJeT 0.76 0.84 0.84 0.89 0.89 0.87 0.80
TJeT 0.74 0.81 0.83 0.87 0.87 0.86 0.76
TJeTAeT 0.76 0.84 0.85 0.89 0.89 0.88 0.80
D.5 Indonesian
D.5.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.97 0.92 0.98 0.94 0.94 0.92 0.88
AeJeT 0.97 0.92 0.97 0.94 0.94 0.92 0.88
AeT 0.97 0.92 0.97 0.94 0.94 0.91 0.89
Greedy 0.97 0.93 0.98 0.95 0.95 0.92 0.89
JeAeT 0.97 0.87 0.97 0.93 0.93 0.91 0.87
JeT 0.97 0.88 0.97 0.93 0.93 0.91 0.87
@@ D.5.1 Precision
TAeT 0.97 0.93 0.98 0.94 0.94 0.91 0.88
TAeTJeT 0.97 0.93 0.98 0.95 0.95 0.91 0.89
TJeT 0.98 0.93 0.99 0.97 0.97 0.94 0.90
TJeTAeT 0.98 0.93 0.99 0.97 0.97 0.94 0.90
D.5.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.74 0.77 0.70 0.86 0.86 0.86 0.87
AeJeT 0.74 0.78 0.70 0.86 0.86 0.86 0.87
AeT 0.74 0.78 0.70 0.86 0.86 0.86 0.88
Greedy 0.78 0.78 0.74 0.87 0.87 0.86 0.90
JeAeT 0.56 0.54 0.57 0.75 0.75 0.68 0.78
JeT 0.55 0.53 0.57 0.75 0.75 0.67 0.80
@@ D.5.2 Recall
TAeT 0.78 0.78 0.72 0.87 0.87 0.85 0.90
TAeTJeT 0.78 0.78 0.73 0.87 0.87 0.85 0.89
TJeT 0.69 0.71 0.75 0.88 0.88 0.82 0.93
TJeTAeT 0.71 0.72 0.74 0.88 0.88 0.83 0.93
D.6 Italian
D.6.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.96 0.79 0.93 0.91 0.91 0.81 0.67
AeJeT 0.96 0.79 0.93 0.91 0.91 0.81 0.65
AeT 0.95 0.79 0.96 0.91 0.91 0.81 0.66
Greedy 0.97 0.82 0.96 0.92 0.92 0.83 0.70
JeAeT 0.96 0.75 0.87 0.86 0.86 0.75 0.54
JeT 0.97 0.76 0.87 0.86 0.86 0.75 0.55
@@ D.6.1 Precision
TAeT 0.96 0.79 0.96 0.91 0.91 0.82 0.67
TAeTJeT 0.97 0.82 0.96 0.92 0.92 0.83 0.68
TJeT 0.98 0.80 0.95 0.93 0.93 0.82 0.67
TJeTAeT 0.97 0.80 0.95 0.92 0.92 0.82 0.69
D.6.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.75 0.88 0.75 0.90 0.90 0.90 0.89
AeJeT 0.75 0.88 0.75 0.90 0.90 0.90 0.88
AeT 0.76 0.89 0.76 0.92 0.92 0.91 0.91
Greedy 0.77 0.89 0.75 0.91 0.91 0.91 0.89
JeAeT 0.72 0.85 0.74 0.88 0.88 0.89 0.88
JeT 0.72 0.84 0.73 0.88 0.88 0.88 0.88
@@ D.6.2 Recall
TAeT 0.77 0.89 0.76 0.92 0.92 0.91 0.92
TAeTJeT 0.77 0.89 0.75 0.91 0.91 0.91 0.90
TJeT 0.75 0.87 0.75 0.90 0.90 0.90 0.90
TJeTAeT 0.76 0.88 0.76 0.90 0.90 0.90 0.90
D.7 Japanese
D.7.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.87 0.82 0.86 0.86 0.86 0.87 0.86
AeJeT 0.87 0.82 0.86 0.86 0.86 0.87 0.85
AeT 0.87 0.83 0.87 0.86 0.86 0.87 0.86
Greedy 0.87 0.84 0.87 0.86 0.86 0.88 0.86
JeAeT 0.86 0.80 0.83 0.84 0.84 0.87 0.85
JeT 0.86 0.80 0.83 0.83 0.83 0.87 0.86
@@ D.7.1 Precision
TAeT 0.87 0.83 0.87 0.86 0.86 0.87 0.86
TAeTJeT 0.87 0.83 0.87 0.86 0.86 0.88 0.86
TJeT 0.87 0.83 0.87 0.84 0.84 0.88 0.86
TJeTAeT 0.87 0.83 0.87 0.85 0.85 0.88 0.86
D.7.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.76 0.52 0.70 0.73 0.73 0.65 0.50
AeJeT 0.76 0.52 0.70 0.73 0.73 0.65 0.50
AeT 0.78 0.55 0.72 0.76 0.76 0.71 0.52
Greedy 0.78 0.55 0.72 0.76 0.76 0.71 0.53
JeAeT 0.76 0.52 0.70 0.73 0.73 0.66 0.49
JeT 0.76 0.52 0.70 0.73 0.73 0.66 0.51
@@ D.7.2 Recall
TAeT 0.78 0.55 0.72 0.76 0.76 0.71 0.52
TAeTJeT 0.78 0.55 0.72 0.76 0.76 0.71 0.52
TJeT 0.78 0.55 0.72 0.76 0.76 0.71 0.52
TJeTAeT 0.78 0.55 0.72 0.76 0.76 0.71 0.52
D.8 Korean
D.8.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.98 0.95 0.98 0.97 0.97 0.98 0.52
AeJeT 0.98 0.95 0.98 0.97 0.97 0.98 0.53
AeT 0.98 0.96 0.98 0.98 0.98 0.98 0.55
Greedy 0.98 0.91 0.98 0.98 0.98 0.98 0.55
JeAeT 0.98 0.94 0.98 0.97 0.97 0.98 0.45
JeT 0.98 0.94 0.98 0.97 0.97 0.98 0.47
@@ D.8.1 Precision
TAeT 0.98 0.96 0.98 0.98 0.98 0.98 0.54
TAeTJeT 0.98 0.91 0.98 0.98 0.98 0.98 0.55
TJeT 0.98 0.90 0.98 0.98 0.98 0.98 0.51
TJeTAeT 0.98 0.91 0.98 0.98 0.98 0.98 0.50
D.8.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.71 0.65 0.71 0.72 0.72 0.73 0.69
AeJeT 0.71 0.65 0.71 0.72 0.72 0.73 0.68
AeT 0.71 0.65 0.70 0.72 0.72 0.73 0.67
Greedy 0.70 0.56 0.71 0.72 0.72 0.73 0.66
JeAeT 0.60 0.55 0.60 0.61 0.61 0.62 0.58
JeT 0.60 0.55 0.60 0.61 0.61 0.62 0.57
@@ D.8.2 Recall
D.9 Portuguese
D.9.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.95 0.92 0.85 0.92 0.92 0.93 0.89
AeJeT 0.95 0.92 0.85 0.92 0.92 0.93 0.89
AeT 0.97 0.92 0.85 0.93 0.93 0.94 0.90
Greedy 0.96 0.92 0.87 0.94 0.94 0.94 0.92
JeAeT 0.91 0.87 0.79 0.86 0.86 0.91 0.86
JeT 0.92 0.89 0.80 0.87 0.87 0.93 0.85
@@ D.9.1 Precision
TAeT 0.96 0.92 0.85 0.93 0.93 0.95 0.89
TAeTJeT 0.96 0.92 0.87 0.94 0.94 0.94 0.91
TJeT 0.97 0.93 0.87 0.95 0.95 0.95 0.92
TJeTAeT 0.96 0.92 0.86 0.94 0.94 0.94 0.90
D.9.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.76 0.81 0.83 0.85 0.85 0.84 0.83
AeJeT 0.76 0.81 0.83 0.85 0.85 0.84 0.82
AeT 0.80 0.84 0.85 0.88 0.88 0.87 0.87
Greedy 0.80 0.86 0.85 0.89 0.89 0.88 0.86
JeAeT 0.68 0.70 0.79 0.76 0.76 0.75 0.77
JeT 0.68 0.71 0.79 0.76 0.76 0.76 0.77
@@ D.9.2 Recall
TAeT 0.81 0.85 0.85 0.88 0.88 0.88 0.86
TAeTJeT 0.80 0.85 0.84 0.88 0.88 0.88 0.88
TJeT 0.77 0.84 0.86 0.88 0.88 0.89 0.88
TJeTAeT 0.77 0.84 0.86 0.88 0.88 0.89 0.89
D.10 Russian
D.10.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.96 0.90 0.96 0.96 0.96 0.96 0.96
AeJeT 0.96 0.90 0.96 0.96 0.96 0.96 0.96
AeT 0.96 0.91 0.96 0.95 0.95 0.95 0.96
Greedy 0.96 0.91 0.96 0.95 0.95 0.95 0.95
JeAeT 0.97 0.87 0.97 0.97 0.97 0.97 0.97
JeT 0.97 0.87 0.97 0.97 0.97 0.97 0.97
@@ D.10.1 Precision
TAeT 0.96 0.91 0.96 0.95 0.95 0.95 0.96
TAeTJeT 0.96 0.91 0.96 0.95 0.95 0.95 0.95
TJeT 0.97 0.91 0.97 0.96 0.96 0.96 0.96
TJeTAeT 0.97 0.92 0.97 0.96 0.96 0.96 0.97
D.10.2 Recall
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.66 0.69 0.69 0.71 0.71 0.68 0.66
AeJeT 0.67 0.70 0.69 0.72 0.72 0.68 0.67
AeT 0.71 0.74 0.74 0.77 0.77 0.73 0.71
Greedy 0.71 0.73 0.70 0.75 0.75 0.74 0.68
JeAeT 0.52 0.58 0.55 0.57 0.57 0.54 0.53
JeT 0.52 0.58 0.55 0.57 0.57 0.54 0.53
@@ D.10.2 Recall
TAeT 0.71 0.74 0.71 0.76 0.76 0.73 0.69
TAeTJeT 0.72 0.73 0.70 0.76 0.76 0.74 0.69
TJeT 0.60 0.62 0.61 0.63 0.63 0.63 0.58
TJeTAeT 0.60 0.62 0.61 0.63 0.63 0.63 0.58
D.11 Spanish
D.11.1 Precision
Corpus CLD2 CLD3
Detect
Language
FastSpell FastText LangID Langdetect
AJeT 0.94 0.92 0.90 0.86 0.86 0.84 0.87
AeJeT 0.94 0.92 0.90 0.86 0.86 0.84 0.87
AeT 0.94 0.93 0.91 0.87 0.87 0.85 0.89
Greedy 0.94 0.92 0.91 0.88 0.88 0.87 0.89
JeAeT 0.95 0.90 0.85 0.78 0.78 0.75 0.81
JeT 0.94 0.90 0.85 0.78 0.78 0.75 0.81

### v4

@@ Appendix B Processing times
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
@@
A 0.19 0.40 3.62 0.15 0.14 5.99 10.15
J 0.07 0.24 2.74 0.08 0.07 3.90 8.94
G 0.13 0.45 3.93 0.16 0.16 4.91 10.18
Appendix C Performance by Language
@@
C.1 Chinese
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
@@ C.1 Chinese
A 0.87 0.75 0.86 0.84 0.84 0.77 0.74
J 0.99 0.76 0.98 0.94 0.94 0.81 0.81
G 0.87 0.75 0.86 0.84 0.84 0.76 0.76
Recall
G 0.76 0.71 0.70 0.77 0.77 0.82 0.37
@@
A 0.75 0.70 0.70 0.77 0.77 0.81 0.39
J 0.70 0.70 0.63 0.70 0.70 0.78 0.29
C.2 English
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.72 0.78 0.70 0.77 0.77 0.79 0.72
@@ C.2 English
A 0.73 0.78 0.70 0.77 0.77 0.78 0.74
J 0.70 0.75 0.69 0.75 0.75 0.76 0.70
Recall
G 0.81 0.71 0.89 0.86 0.86 0.81 0.76
@@
A 0.78 0.69 0.88 0.86 0.86 0.80 0.73
J 0.85 0.73 0.93 0.91 0.91 0.83 0.79
C.3 French
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.90 0.84 0.92 0.86 0.86 0.82 0.76
@@ C.3 French
A 0.88 0.84 0.92 0.86 0.86 0.81 0.76
J 0.96 0.89 0.95 0.92 0.92 0.86 0.82
Recall
G 0.77 0.85 0.78 0.91 0.91 0.89 0.89
@@
A 0.78 0.85 0.77 0.90 0.90 0.89 0.89
J 0.75 0.84 0.79 0.90 0.90 0.89 0.90
C.4 German
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.88 0.83 0.89 0.87 0.87 0.81 0.86
@@ C.4 German
A 0.88 0.83 0.89 0.87 0.87 0.80 0.85
J 0.93 0.86 0.92 0.91 0.91 0.83 0.89
Recall
G 0.76 0.84 0.85 0.89 0.89 0.88 0.79
@@
A 0.76 0.82 0.84 0.89 0.89 0.87 0.79
J 0.74 0.81 0.83 0.87 0.87 0.86 0.76
C.5 Indonesian
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.97 0.93 0.98 0.95 0.95 0.92 0.89
@@ C.5 Indonesian
A 0.97 0.93 0.98 0.94 0.94 0.91 0.88
J 0.98 0.93 0.99 0.97 0.97 0.94 0.90
Recall
G 0.78 0.78 0.74 0.87 0.87 0.86 0.90
@@
A 0.78 0.78 0.72 0.87 0.87 0.85 0.90
J 0.69 0.71 0.75 0.88 0.88 0.82 0.93
C.6 Italian
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.97 0.82 0.96 0.92 0.92 0.83 0.70
@@ C.6 Italian
A 0.96 0.79 0.96 0.91 0.91 0.82 0.67
J 0.98 0.80 0.95 0.93 0.93 0.82 0.67
Recall
G 0.77 0.89 0.75 0.91 0.91 0.91 0.89
@@
A 0.77 0.89 0.76 0.92 0.92 0.91 0.92
J 0.75 0.87 0.75 0.90 0.90 0.90 0.90
C.7 Japanese
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.87 0.84 0.87 0.86 0.86 0.88 0.86
@@ C.7 Japanese
A 0.87 0.83 0.87 0.86 0.86 0.87 0.86
J 0.87 0.83 0.87 0.84 0.84 0.88 0.86
Recall
G 0.78 0.55 0.72 0.76 0.76 0.71 0.53
@@
A 0.78 0.55 0.72 0.76 0.76 0.71 0.52
J 0.78 0.55 0.72 0.76 0.76 0.71 0.52
C.8 Korean
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.98 0.91 0.98 0.98 0.98 0.98 0.55
@@ C.8 Korean
A 0.98 0.96 0.98 0.98 0.98 0.98 0.54
J 0.98 0.90 0.98 0.98 0.98 0.98 0.51
Recall
G 0.70 0.56 0.71 0.72 0.72 0.73 0.66
@@
C.9 Portuguese
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.96 0.92 0.87 0.94 0.94 0.94 0.92
@@ C.9 Portuguese
A 0.96 0.92 0.85 0.93 0.93 0.95 0.89
J 0.97 0.93 0.87 0.95 0.95 0.95 0.92
Recall
G 0.80 0.86 0.85 0.89 0.89 0.88 0.86
@@
A 0.81 0.85 0.85 0.88 0.88 0.88 0.86
J 0.77 0.84 0.86 0.88 0.88 0.89 0.88
C.10 Russian
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.96 0.91 0.96 0.95 0.95 0.95 0.95
@@ C.10 Russian
A 0.96 0.91 0.96 0.95 0.95 0.95 0.96
J 0.97 0.91 0.97 0.96 0.96 0.96 0.96
Recall
G 0.71 0.73 0.70 0.75 0.75 0.74 0.68
@@
A 0.71 0.74 0.71 0.76 0.76 0.73 0.69
J 0.60 0.62 0.61 0.63 0.63 0.63 0.58
C.11 Spanish
Corpus
CLD2
CLD3
Detect Language
FastSpell
FastText
LangID
Langdetect
Precision
G 0.94 0.92 0.91 0.88 0.88 0.87 0.89